OpenStudy (anonymous):

what is the length of the hypotenuse of a right angle with the legs of 9cm and 12 cm. Fan and medal

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

need i say more? :)

OpenStudy (anonymous):

Using proportions of corresponding sides we have CD/9 = 12/15 CD = 9*12/15 CD = 36/5 Answer: the line is 7.2 cm

OpenStudy (anonymous):

yes I WILL GIVE YOU MORE MEDALS!!!!!!

OpenStudy (anonymous):

whats answer choices?

OpenStudy (anonymous):

okay so 9^2+12^2=? @skullcrusher101

OpenStudy (anonymous):

225

OpenStudy (anonymous):

oops 15

OpenStudy (anonymous):

YOU ARE TO LATE @atreyu6s3x6

OpenStudy (anonymous):

225 is correct so we get: 9^2+12^2=225^2 so we square root to eliminate the square: \[\sqrt{225}\]=?

OpenStudy (anonymous):

sorry @skullcrusher101 dont crush me... :( haha but rally my bad, it takes time to type this out

OpenStudy (anonymous):

its good you want a fan?

OpenStudy (anonymous):

^^^^^^^^^^ the answer is 15

OpenStudy (anonymous):

sure and yes it is :D good work

OpenStudy (anonymous):

get my next question

OpenStudy (anonymous):

it hasnt popped up..@skullcrusher101

OpenStudy (anonymous):

post it here

OpenStudy (anonymous):

I will psot everything here. The

OpenStudy (anonymous):

Got to go